OpenStudy (anonymous):

Suppose there is a rope tied around the equator of the Earth, and the equator is a perfect circle. If I were to add 1 meter to the rope (assuming it spreads out evenly) what would the distance between the Earth and the rope? The radius of the Earth is 6400 km.

Parth (parthkohli):

Oh my, my. I've had this question asked by my cousin.

Parth (parthkohli):

\[l = 2 \pi r = 12800 \pi ~km\]\[\text{New length, }l_0 = 2\pi r + 1 = 12800 \pi + 1 = 2\pi R_0\]Where \(R_0\) is the new radius after adding 1 metre. Now, if you think about it carefully, the distance between the Earth and the rope is actually the difference in radii, right? So you need to find \(R_0 - r\)

Parth (parthkohli):

\[R_0 = \dfrac{12800\pi + 1}{2\pi} = 6400 + \frac{1}{2\pi}\]\[r = 6400\]Therefore, the distance between the rope and the Earth\[R_0 - r = \dfrac{1}{2\pi }\]

OpenStudy (anonymous):

ParthKholi do you have phd in math

OpenStudy (anonymous):

What does the R0 and L0 stand for?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

I've written in my answer that \(R_0\) is the new radius of the rope after 1 m is added and \(l_0\) is the new length.